Twice the square of a number is 10 more than the number. What is the number?
Answer by mananth(16949)

number =x
2x^2=x+10
2x^2-x-10=0
2x^2-5x+4x-10=0
x(2x-5)+2(2x-5)=0
(2x-5)(x+2)=0
x=5/2 OR x=-2